# Representations of knot groups in $\textup{AGL}_{1}(\mathbb{C})$ and Alexander invariants

### v1

@@ 5.1. The tangle category
We outline the main definitions and properties of the tangle category; for a detailed exposition we refer to [26]. For any integer $n\geq 1$, let $[n]$ denote the set $\{1,\ldots,n\}$ and set $[0]=\emptyset$. A tangle $T$ is the union of a finite number of disjoint oriented polygonal arcs in $\mathbb{R}^{2}\times[0,1]$ such that its boundary satisfies the condition
@@
  • Objects: finite sequences of $\pm$ signs, possibly empty. Given any sequence $\epsilon=(\epsilon_{1},\ldots,\epsilon_{n})$ with $\epsilon_{i}=\pm$, we shall write $-\epsilon=(-\epsilon_{1},\ldots,-\epsilon_{n})$ for the sequence where all signs have been flipped, and $|\epsilon|=n$ for its length. Equivalently, an object $\epsilon$ can be seen as the complementary surface $\epsilon^{c}=\mathbb{R}^{2}-\left([|\epsilon|]\times\{0\}\right)$ with a choice of orientations for the normal bundle around the punctures, according to the signs $\epsilon_{i}$.
@@
    Given any object $\epsilon$ of ${\mathcal{T}}$, the identity morphism $\operatorname{Id}_{\epsilon}$ is the tangle formed by the union of intervals $[|\epsilon|]\times\{0\}\times[0,1]$, where the orientation of the intervals is determined so that $s(\operatorname{Id}_{\epsilon})=b(\operatorname{Id}_{\epsilon})=\epsilon$. In particular, we will write $\operatorname{Id}_{+}=\,\uparrow$ and $\operatorname{Id}_{-}=\,\downarrow\,$.
@@
Tangles can be represented using tangle diagrams, which are tangle projections to ×$\mathbb{R}\times[0,1]$ so that every crossing is simple. As it occurs for links and diagrams, two tangle diagrams represent isotopic tangles if one is obtained from the other by a sequence of Reidemeister transformations and isotopies of diagrams.
@@
,$,\curvearrowright,\curvearrowleft,X_{+},X_{-},$ , (20)
subject to a collection of relations, detailed in [26, Theorem XII.2.2], which arise from isotopies of diagrams and Reidemeister transformations.
@@ 5.2. The category $\operatorname{Span}(\operatorname{\mathbf{Vect}}_{{\mathbb{C}}})$
    ${Z\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$$\scriptstyle{g}$${X}$${Y}$ (21)
@@
    ${Z\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$$\scriptstyle{g}$$\scriptstyle{\varphi}$${X}$${Y}$${Z^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{\prime}}$$\scriptstyle{g^{\prime}}$
@@
[…]
@@
${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{Id}_{X}}$$\scriptstyle{f}$${X}$${Y}$
@@
${X=X\times_{Y}Y\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{Id}_{X}}$$\scriptstyle{f}$${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{Id}_{X}}$$\scriptstyle{f}$${Y\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{Id}_{Y}}$$\scriptstyle{g}$${X}$${Y}$${Z}$ (22)
@@
${Z\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$$\scriptstyle{g}$$\scriptstyle{f}$${X}$${Y}$${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{Id}_{X}}$$\scriptstyle{g\circ f^{-1}}$
@@
[…]
Notice that $\oplus$ is both the product and coproduct in $\operatorname{\mathbf{Vect}}_{\mathbb{C}}$, so $(\operatorname{\mathbf{Vect}}_{\mathbb{C}},\oplus)$ is a cartesian and cocartesian category. The monoidal unit of this structure is the zero vector space 00.
@@ 5.3. The Topological Quantum Field Theory
    where $Q_{\epsilon}$ is the quandle associated to the oriented codimension two submanifold $[|\epsilon|]\times\{0\}\subset{\mathbb{R}}^{2}$ given by the points of $\epsilon$ with the orientations given by the signs $\epsilon_{i}$. Its quandle $Q_{\epsilon}$ is the free quandle generated by these points, and $\operatorname{As}(Q(\epsilon^{c}))=\pi_{1}(\epsilon^{c})$, where $\epsilon^{c}={\mathbb{R}}^{2}-\left([|\epsilon|]\times\{0\}\right)$ is the plane punctured at the points of $\epsilon$. It follows that $\pi_{1}(\epsilon^{c})=\mathbf{F}_{|\epsilon|}$, the free group in $|\epsilon|$ generators.
@@
Given any object $\epsilon$ of ${\mathcal{T}}$ of length $n=|\epsilon|$, the identity tangle $\operatorname{Id}_{\epsilon}$ is given by the union of intervals $[n]\times\{0\}\times[0,1]\subset\mathbb{R}^{2}\times[0,1]$, that is, $n$ intervals with the same orientation. Its fundamental group is free in $n$ generators, so ${\mathcal{R}}_{{\textup{AQ}_{t}}}(\operatorname{Id}_{\epsilon}^{c})={\mathbb{C}}^{n}$ and the restriction maps to both boundaries ${\mathcal{R}}_{{\textup{AQ}_{t}}}(\epsilon^{c})={\mathbb{C}}^{n}$ are the identity. Thus, the associated span ${\mathcal{A}}_{t}(\operatorname{Id}_{\epsilon})=({\mathbb{C}}^{n},\operatorname{Id}_{{\mathbb{C}}^{n}},\operatorname{Id}_{{\mathbb{C}}^{n}})$ is the identity span.
@@
[…]
@@
[…]
@@ 5.4. TQFT for singular vector bundles
${E^{\prime\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(f,\tilde{f})}$$\scriptstyle{(g,\tilde{g})}$${E}$${E^{\prime}}$
@@
    where $\Gamma_{\epsilon}=\pi_{1}(\epsilon^{c})$ is the fundamental group of the complement of the points of $[|\epsilon|]\times\{0\}\subset{\mathbb{R}}^{2}$. Under the map ${\textup{AGL}_{1}}\to{\mathbb{C}}^{*}$, we see ${\mathcal{R}}_{{\textup{AGL}_{1}}}(\epsilon)$ as a singular vector bundle over ${\mathcal{R}}_{{\mathbb{C}}^{*}}(\epsilon)$ as in Section 3.4.
@@
[…]
@@ 6.1. Action on the generators of ${\mathcal{T}}$
  • Objects. Given an object $\epsilon$ of ${\mathcal{T}}$ of length $n=|\epsilon|$, we have $\Gamma_{\epsilon}=\pi_{1}(\epsilon^{c})=\mathbf{F}_{n}$, with generators looping counterclockwise (as seen ‘from above’) for positive oriented points $+$, and loop clockwise for negative oriented points --, as shown in Figure 4.
@@
Incoming $\begin{array}[]{rcl}\textrm{Incoming boundary:}&\mathbf{F}_{2}=\pi_{1}({(+,+)^{c}})\to\mathbf{F}_{2}=\pi_{1}(X_{+}^{c}),&(\beta_{1},\beta_{2})\mapsto(\alpha_{1},\alpha_{2}),\\ \textrm{Outgoing boundary:}&\mathbf{F}_{2}=\pi_{1}({(+,+)^{c}})\to\mathbf{F}_{2}=\pi_{1}(X_{+}^{c}),&(\beta_{1},\beta_{2})\mapsto(\alpha_{1}\alpha_{2}\alpha_{1}^{-1},\alpha_{1})=(\alpha_{1}\triangleright\alpha_{2},\alpha_{1}).\end{array}$
@@
${{\mathbb{C}}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{Id}}$$\scriptstyle{f_{+}}$${{\mathbb{C}}^{2}}$${{\mathbb{C}}^{2}}$
@@
${{\mathbb{C}}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{Id}}$$\scriptstyle{f_{-}}$${{\mathbb{C}}^{2}}$${{\mathbb{C}}^{2}}$
@@
[…]
@@
[…]
@@ 6.2. The braid subcategory
The category ${\mathcal{T}}$ of tangles admits a distinguished subcategory ${\mathcal{B}}$, called the category of braids. It contains the objects of the form (+,,+)(+$\ldots$,+)( + , … , + ) of ${\mathcal{T}}$, that we will simply denote by its length $n$. As morphisms, we only allow tangles always moving upwards i.e. those tangles whose tangent line is never horizontal, as shown in Figure 9.
@@
${{\mathcal{T}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\mathcal{A}}_{t}}$${\operatorname{Span}(\operatorname{\mathbf{Vect}}_{\mathbb{C}})}$${{\mathcal{B}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\mathcal{A}}_{t}|_{{\mathcal{B}}}}$${\operatorname{\mathbf{Vect}}_{\mathbb{C}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@
[…]
@@
[…]
@@
${\mathbb{C}}^{n}\oplus_{{\mathbb{C}}^{2n}}{\mathbb{C}}^{n}$ =$=\left\{(v_{1},v_{2})\in{\mathbb{C}}^{n}\oplus{\mathbb{C}}^{n}\,\,|\,\,(f\oplus\operatorname{Id}_{n})\circ\Delta(v_{1})=\Delta(v_{2})\right\}$
=$=\left\{(v_{1},v_{2})\in{\mathbb{C}}^{n}\oplus{\mathbb{C}}^{n}\,\,|\,\,f(v_{1})=v_{2},\,v_{1}=v_{2}\right\}=\left\{v\in{\mathbb{C}}^{n}\,\,|\,\,f(v)=v\right\}=\operatorname{Fix}(f),$
@@
In Corollary 6.7, any other column or row can be removed to compute the determinant. However, since the Alexander polynomial is only determined up to a factor ±$\pm t^{k}$, it turns out that the resulting polynomial will differ by one of these factors. Making the choice of removing the first row and column leads to a polynomial with positive constant term, which coincides with the usual normalization of the Alexander polynomial.

### v2

@@ 5.1. The tangle category
We outline the main definitions and properties of the tangle category; for a detailed exposition we refer to [33]. For any integer $n\geq 1$, let $[n]$ denote the set $\{1,\ldots,n\}$ and set $0]=\emptyset$. A tangle $T$ is the union of a finite number of disjoint oriented polygonal arcs in $\mathbb{R}^{2}\times[0,1]$ such that its boundary satisfies the condition
@@
  • Objects: finite sequences of $\pm$ signs, possibly empty. Given any sequence $\epsilon=(\epsilon_{1},\ldots,\epsilon_{n})$ with $\epsilon_{i}=\pm$, we shall write $\epsilon=(-\epsilon_{1},\ldots,-\epsilon_{n})$ for the sequence where all signs have been flipped, and $|\epsilon|=n$ for its length. Equivalently, an object $\epsilon$ can be seen as the complementary surface $\epsilon^{c}=\mathbb{R}^{2}-\left([|\epsilon|]\times\{0\}\right)$ with a choice of orientations for the normal bundle around the punctures, according to the signs $\epsilon_{i}$.
@@
    Given any object $\epsilon$ of $\mathcal{T}$}, the identity morphism $\operatorname{Id}_{\epsilon}$ is the tangle formed by the union of intervals $|\epsilon|]\times\{0\}\times[0,1]$, where the orientation of the intervals is determined so that $s(\operatorname{Id}_{\epsilon})=b(\operatorname{Id}_{\epsilon})=\epsilon$. In particular, we will write $\operatorname{Id}_{+}=\,\uparrow$ and $\operatorname{Id}_{-}=\,\downarrow\,$.
@@
Tangles can be represented using tangle diagrams, which are tangle projections to $\mathbb{R}\times[0,1]$ so that every crossing is simple. As it occurs for links and diagrams, two tangle diagrams represent isotopic tangles if one is obtained from the other by a sequence of Reidemeister transformations and isotopies of diagrams.
@@
[…]
subject to a collection of relations, detailed in [33, Theorem XII.2.2], which arise from isotopies of diagrams and Reidemeister transformations.
@@ 5.2. The category $\operatorname{Span}(\operatorname{\mathbf{Vect}}_{{\mathbb{C}}})$
[…]
@@
[…]
@@
    ${Z_{1}\times_{Y}Z_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{1}}$$\scriptstyle{\pi_{2}}$${Z_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{1}}$$\scriptstyle{g_{1}}$${Z_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{2}}$$\scriptstyle{g_{2}}$${X}$${Y}$${Z}$
@@
[…]
@@
[…]
@@
[…]
@@
${Z\oplus Z^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f\oplus f^{\prime}}$$\scriptstyle{g\oplus g^{\prime}}$${X\oplus X^{\prime}}$${Y\oplus Y^{\prime}}$
Notice that $\oplus$ is both the product and coproduct in $\operatorname{\mathbf{Vect}}_{\mathbb{C}}$, so $(\operatorname{\mathbf{Vect}}_{\mathbb{C}},\oplus)$ is a cartesian and cocartesian category. The monoidal unit of this structure is the zero vector space 0.
@@ 5.3. The Topological Quantum Field Theory
    where $Q_{\epsilon}$ is the quandle associated to the oriented codimension two submanifold $|\epsilon|]\times\{0\}\subset{\mathbb{R}}^{2}$ given by the points of $\epsilon$ with the orientations given by the signs $\epsilon_{i}$. Its quandle $Q_{\epsilon}$ is the free quandle generated by these points, and $\operatorname{As}(Q(\epsilon^{c}))=\pi_{1}(\epsilon^{c})$, where $\epsilon^{c}={\mathbb{R}}^{2}-\left([|\epsilon|]\times\{0\}\right)$ is the plane punctured at the points of $\epsilon$. It follows that $\pi_{1}(\epsilon^{c})=\mathbf{F}_{|\epsilon|}$, the free group in $|\epsilon|$ generators.
@@
Given any object $\epsilon$ of $\mathcal{T}$} of length $n=|\epsilon|$, the identity tangle $\operatorname{Id}_{\epsilon}$ is given by the union of intervals $n]\times\{0\}\times[0,1]\subset\mathbb{R}^{2}\times[0,1]$, that is, $n$ intervals with the same orientation. Its fundamental group is free in $n$ generators, so ${\mathcal{R}}_{{\textup{AQ}_{t}}}(\operatorname{Id}_{\epsilon}^{c})={\mathbb{C}}^{n}$ and the restriction maps to both boundaries ${\mathcal{R}}_{{\textup{AQ}_{t}}}(\epsilon^{c})={\mathbb{C}}^{n}$ are the identity. Thus, the associated span ${\mathcal{A}}_{t}(\operatorname{Id}_{\epsilon})=({\mathbb{C}}^{n},\operatorname{Id}_{{\mathbb{C}}^{n}},\operatorname{Id}_{{\mathbb{C}}^{n}})$ is the identity span.
@@
${{\mathcal{R}}_{{\textup{AQ}_{t}}}(T^{\prime}\circ T)\ \ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{R}}_{{\textup{AQ}_{t}}}(T)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(i_{1})_{*}}$${{\mathcal{R}}_{{\textup{AQ}_{t}}}(T^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(i_{2})_{*}}$${{\mathcal{R}}_{{\textup{AQ}_{t}}}(\epsilon_{1})}$${{\mathcal{R}}_{{\textup{AQ}_{t}}}(\epsilon_{2})}$${{\mathcal{R}}_{{\textup{AQ}_{t}}}(\epsilon_{3})}$
@@
${{\mathcal{R}}_{{\textup{AQ}_{t}}}(T)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${0}$
@@ 5.4. TQFT for singular vector bundles
[…]
@@
    where $\Gamma_{\epsilon}=\pi_{1}(\epsilon^{c})$ is the fundamental group of the complement of the points of $|\epsilon|]\times\{0\}\subset{\mathbb{R}}^{2}$. Under the map $\textup{AGL}_{1}$$\to{\mathbb{C}}^{*}$, we see ${\mathcal{R}}_{{\textup{AGL}_{1}}}(\epsilon)$ as a singular vector bundle over ${\mathcal{R}}_{{\mathbb{C}}^{*}}(\epsilon)$ as in Section 3.4.
@@
${{\mathcal{R}}_{{\textup{AGL}_{1}}}(T)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${0}$
@@ 6.1. Action on the generators of $\mathcal{T}$}
  • Objects. Given an object $\epsilon$ of $\mathcal{T}$} of length $n=|\epsilon|$, we have $\Gamma_{\epsilon}=\pi_{1}(\epsilon^{c})=\mathbf{F}_{n}$, with generators looping counterclockwise (as seen ‘from above’) for positive oriented points ++, and loop clockwise for negative oriented points -, as shown in Figure 4.
@@
$\begin{array}[]{rcl}\textrm{Incoming boundary:}&\mathbf{F}_{2}=\pi_{1}({(+,+)^{c}})\to\mathbf{F}_{2}=\pi_{1}(X_{+}^{c}),&(\beta_{1},\beta_{2})\mapsto(\alpha_{1},\alpha_{2}),\\ \textrm{Outgoing boundary:}&\mathbf{F}_{2}=\pi_{1}({(+,+)^{c}})\to\mathbf{F}_{2}=\pi_{1}(X_{+}^{c}),&(\beta_{1},\beta_{2})\mapsto(\alpha_{1}\alpha_{2}\alpha_{1}^{-1},\alpha_{1})=(\alpha_{1}\triangleright\alpha_{2},\alpha_{1}).\end{array}$
@@
[…]
@@
[…]
@@
${{\mathcal{R}}_{{\textup{AQ}_{t}}}(\epsilon)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Delta}$${{\mathcal{R}}_{{\textup{AQ}_{t}}}(\epsilon)^{\oplus 2}}$${0}$
@@
${{\mathcal{R}}_{{\textup{AQ}_{t}}}(\epsilon)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Delta}$${0}$${{\mathcal{R}}_{{\textup{AQ}_{t}}}(\epsilon)^{\oplus 2}}$
@@ 6.2. The braid subcategory
The category $\mathcal{T}$} of tangles admits a distinguished subcategory $\mathcal{B}$}, called the category of braids. It contains the objects of the form $(+,\ldots,+)$ of $\mathcal{T}$}, that we will simply denote by its length $n$. As morphisms, we only allow tangles always moving upwards i.e. those tangles whose tangent line is never horizontal, as shown in Figure 9.
@@
${{\mathcal{T}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\mathcal{A}}_{t}}$${\operatorname{Span}(\operatorname{\mathbf{Vect}}_{\mathbb{C}})}\textstyle{{\mathcal{B}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\mathcal{A}}_{t}|_{{\mathcal{B}}}}$${\operatorname{\mathbf{Vect}}_{\mathbb{C}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@
${\operatorname{Fix}(f)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${0}$
@@
${{\mathbb{C}}^{2n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{Id}_{2n}}$$\scriptstyle{f\oplus\operatorname{Id}_{n}}$${{\mathbb{C}}^{2n}}$${{\mathbb{C}}^{2n}}$
@@
${\mathbb{C}}^{n}\oplus_{{\mathbb{C}}^{2n}}{\mathbb{C}}^{n}$ $=\left\{(v_{1},v_{2})\in{\mathbb{C}}^{n}\oplus{\mathbb{C}}^{n}\,\,|\,\,(f\oplus\operatorname{Id}_{n})\circ\Delta(v_{1})=\Delta(v_{2})\right\}$
$=\left\{(v_{1},v_{2})\in{\mathbb{C}}^{n}\oplus{\mathbb{C}}^{n}\,\,|\,\,f(v_{1})=v_{2},\,v_{1}=v_{2}\right\}=\left\{v\in{\mathbb{C}}^{n}\,\,|\,\,f(v)=v\right\}=\operatorname{Fix}(f),$
@@
In Corollary 6.7, any other column or row can be removed to compute the determinant. However, since the Alexander polynomial is only determined up to a factor $\pm t^{k}$, it turns out that the resulting polynomial will differ by one of these factors. Making the choice of removing the first row and column leads to a polynomial with positive constant term, which coincides with the usual normalization of the Alexander polynomial.